Problem 5-17 (LO 5-3, 5-4, 5-5)

On January 1, 2017, Corgan Company acquired 80 percent of the outstanding voting stock of Smashing, Inc., for a total of $1,440,000 in cash and other consideration. At the acquisition date, Smashing had common stock of $870,000, retained earnings of $420,000, and a noncontrolling interest fair value of $360,000. Corgan attributed the excess of fair value over Smashing's book value to various covenants with a 20-year remaining life. Corgan uses the equity method to account for its investment in Smashing.

During the next two years, Smashing reported the following:

Net Income Dividends Declared Inventory Purchases from Corgan
2017 $ 320,000 $ 52,000 $ 270,000
2018 300,000 62,000 290,000

Corgan sells inventory to Smashing using a 60 percent markup on cost. At the end of 2017 and 2018, 40 percent of the current year purchases remain in Smashing's inventory.

  1. Compute the equity method balance in Corgan's Investment in Smashing, Inc., account as of December 31, 2018.
  2. Prepare the worksheet adjustments for the December 31, 2018, consolidation of Corgan and Smashing.
Answer #1

Pig - Answer: © Equity method Balance: Acquisition Aualysis Consideration transferred by Corgan =$14, 40,000 farv value of noPig ® Amortization per year - initial value / lite fyers] $ 5,10,000 = 25,500 - 20 Amortization per year = $25,500 om POMY capag ☺ Year 2018 ice CANTAN Inventory Purchase from Corgan - $ 2.90,000 6044 Profit on Cost cost Price ($ 2,90,000/160-%) = $Pge 40500 (43500) 216600 Deffered profit in intra company transfer 10 2011 Deffered profit in infra company toansfer V 2018 2P.g. Date | Gredit Debit 40,500 40,500 8,70,000 6,88,000 - particular Investment in Samshing cost of goods sold Common stock